A coil with inductance of 3.0 H and a resistance of 34.0 ? is suddenly connected to a resistanceless battery with ? = 100.0 V, by closing the switch.

What is the equilibrium current?

How much energy is stored in the magnetic field when this current exists in the coil?
Answer #1

equilibrium current = 100/34=2.9411764 A

energy stored= 1/2*(LI^2)= 12*3*(100/34)^2=12.97